OpenStudy (mathisfun13):

find the sum of this infinite geometric series: 100+60+36+...

OpenStudy (anonymous):

It's fairly obvious to see that the common ratio is:\[\bf Common \ ratio = \frac{60}{100}=\frac{36}{60}=\frac{3}{5}\]So the geometric series is given by ('a' is the first term; 'r' is the common ratio):\[\bf ar^{n-1}=100\left( \frac{ 3 }{ 5 } \right)^{n-1}\]Since this geometric series is convergent, i.e. \(\bf |r| < 1\), the sum of the series is given by:\[\bf S_n=\frac{ a }{ 1-r }\]Can you evaluate the sum? @mathisfun13

OpenStudy (mathisfun13):

uhmm i think so

OpenStudy (mathisfun13):

is the r my 3/5?

OpenStudy (mathisfun13):

....

OpenStudy (zzr0ck3r):

yes

OpenStudy (mathisfun13):

so whats my Sn

terenzreignz (terenzreignz):

I think the formula has already been neatly provided :) \[\Large S_n=\frac{a}{1-r}\] Where r is the common ratio and a is the first term in the series :)

OpenStudy (mathisfun13):

ooh okay so 100 would be my a

OpenStudy (zzr0ck3r):

correct

terenzreignz (terenzreignz):

Yup :)

OpenStudy (mathisfun13):

i got 250

OpenStudy (anonymous):

Correct.